OpenStudy (anonymous):

please help me round the number to the given place value. 1.044; hundredths a. 1.5 b. 1.4 c. 1.05 d. 1.04

OpenStudy (anonymous):

D

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

any more questions? =)

OpenStudy (anonymous):

yes estimate the value of the expression by rounding each number to the given place value. 2968+9912; thousands a. 14,000 b. 13,000 c. 12,000 d. 7000

OpenStudy (anonymous):

hello anybody here

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[2968 + 9912 = 1[2]880\] Thousands place is in parenthesis[that doesn't change anything], how would you round that?

OpenStudy (anonymous):

2968 rounds up to 3000 and 9912 rounds to 10000

OpenStudy (anonymous):

b. 13000

OpenStudy (anonymous):

because you are adding 29 68 and 9912, if you round, you get 3000 and 10000, add them = 13000

OpenStudy (anonymous):

*2912

OpenStudy (anonymous):

*2968 i mean

OpenStudy (anonymous):

estimate the value of the expression by rounding to the given place value. 58/5; round 58 to the nearest 10 a. 10 b. 12 c. 14 d. 20

OpenStudy (anonymous):

58 / 5 = 11.6. Round that to 12. B But it says to round 58 to the nearest ten. So 60 / 5 = 12. so its still B

OpenStudy (anonymous):

so your answer is B

OpenStudy (anonymous):

what are the lower bound estimate and the upper bound estimate found by rounding to the greatest place. 672-324 a. lower bound 300 upper bound 500 b. lower bound 200 upper bound 300 c. lower bound 200 upper bound 400 d. lower bound 300 upper bound 400

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

So the largest the difference could be is if you round to 700 - 300...which is 400. The smallest the difference could be is if you round to 600 - 400... which is 200.

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

1 more question

OpenStudy (anonymous):

k

OpenStudy (anonymous):

which number rounds up when rounded to the nearest hundredth. a. 7.246 b. 9.394 c. 10.002 d. 12.071

OpenStudy (anonymous):

[A = 7.246 \approx7.250\]

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

michel and susan have a combined age of 17 years. michel is 5 years older than susan. how old is michel a.5 b.6 c.11 d.22

OpenStudy (anonymous):

any body here

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

C

OpenStudy (anonymous):

sorry for all the late replies, im pre occupied with organizing my google calendar